# MovieLLM: Enhancing Long Video Understanding with AI-Generated Movies

### v1

@@ Abstract
The development of multimodal models has marked a significant step forward in how machines understand videos. These models have shown promise in analyzing short video clips. However, when it comes to longer formats like movies, they often fall short. The main hurdles are the lack of high-quality, diverse video data and the intensive work required to collect or annotate such data. In the face of these challenges, we propose MovieLLM, a novel framework designed to create synthetic, high-quality data for long videos. This framework leverages the power of GPT-4 and text-to-image models to generate detailed scripts and corresponding visuals. Our approach stands out for its flexibility and scalability, making it a superior alternative to traditional data collection methods. Our extensive experiments validate that the data produced by MovieLLM significantly improves the performance of multimodal models in understanding complex video narratives, overcoming the limitations of existing datasets regarding scarcity and bias.
@@
[…]
Figure 1: Examples of generated long video instruction data. We use GPT-4 and guided text-to-image generation models to generate consistent key frames of move-level video with reasonable lines and corresponding question-answer pairs. These data are used to train multimodal large language models on video understanding.
11footnotetext: The first three authors contributed equally to this work.22footnotetext: Project Leader.33footnotetext: Corresponding Author.
@@ 1 Introduction
In recent years, the field of artificial intelligence has witnessed significant progress in the development of large language models (LLMs) such as GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023). These models have shown exceptional ability in generating coherent and contextually relevant text. Further extending the capabilities of LLMs into the multimodal domain, vision language models (VLMs) like LLaVA (Liu et al., 2023a) have demonstrated remarkable proficiency in tasks that require an understanding of both visual and linguistic information, such as image captioning and visual question-answering (QA). The evolution of VLMs has now begun to include the dynamic realm of video understanding, with models like Video-LLaMA (Zhang et al., 2023) and VideoChat (Li et al., 2023b) highlighting the potential of VLMs in processing and interpreting video content.
However, a notable gap remains in the ability of these models to understand long-duration videos, such as full-length movies. This limitation is primarily attributed to the lack of extensive long video datasets necessary for tuning these models. The LLaMA-VID (Li et al., 2023c) initiative represents a step forward in this direction by creating long-video tuning datasets derived from MovieNet (Huang et al., 2020), indicating that finetuning LLMs on multimodal long video datasets can enhance their movie-level video comprehension. Nevertheless, constructing such benchmarks from existing datasets is both cost-intensive and challenging, particularly due to the manual effort required to collect or annotate additional information like movie dialogues. Moreover, despite showing promise, the QA dataset built upon MovieNet confronts several hurdles. These include the laborious nature of manual data collection and the limited diversity in content, which can impede the generalization capabilities of models due to inherent dataset biases.
Motivated by these challenges, our paper introduces a novel and flexible approach for generating comprehensive datasets for long video instruction tuning. Our methodology leverages the linguistic prowess of GPT-4 and the generative power of stable diffusion models, offering a unique solution to the problem of dataset diversity and richness. Our pipeline comprises three primary stages:
@@
2) Style Immobilization Process. By adeptly employing textual inversion (Gal et al., 2022), we immobilize the style descriptions generated from the script onto the latent space of the diffusion model. This approach guides the model to generate scenes in a fixed style while maintaining diversity under a unified aesthetic.
3) Video instruction data generation. By integrating the powerful generative capabilities of GPT-4 with the developed style-guided diffusion model, we produce style-consistent key frames and corresponding QA pairs, resulting in a comprehensive instruction tuning corpus, combining the visual data with QA pairs.
@@
  • We develop a novel pipeline for generating movie-level video instruction tuning datasets by combining GPT-4 and diffusion models.
@@
  • Based on a real movie dataset, we propose a benchmark for evaluating long video comprehension capabilities. Experiments on the benchmark show the effectiveness of the proposed method, significantly outperforming the baseline.
@@
2 Preliminary
To generate key frames with coherent scenes, we employed the technique, textual inversion (Gal et al., 2022). This preliminary section provides an overview of the textual inversion to lay the groundwork for a detailed exposition of our pipeline.
Textual Inversion
This technique aims to enable language-guided image generation of new, user-specified concepts in a Text-to-Image (T2I) model. To achieve this objective, two main steps are required: 1) text-embedding, this process first converts a string incorporating placeholder words into tokens. Then, these tokens are converted to a continuous vector representation. 2) inversion process, the embedding is then transformed into a single conditioning code that guides the generative model. The embedding vector associated with the so-called pseudo-word will be optimized using a reconstruction objective. As such, the embedding is motivated to capture visual details unique to the concept. The learned embedding is then used as a condition to guide the T2I model to generate customized images.
3 Method
This section elucidates our novel approach for generating movie-level video instruction tuning data. Our method compromises three interrelated steps: movie plot generation, style immobilization process, and video instruction data generation. The complete process is depicted in Fig. 2. The details of these steps are provided in the following subsections.
Figure 3: Illustration of the story expansion strategy. To obtain frame descriptions of a movie-level video and maximally mitigate the forgetting issue, we adopt a three-level story expansion strategy. The strategy can be divided into three levels: (a) epoch chapter generation, where each chapter represents a significant period in story development. (b) narrative threads generation, where each thread refines each major chapter into more specific narrative sub-stories. (c) frame description generation, where each frame description elaborates on these narrative sub-stories, depicting the critical visual frames.
3.1 Movie Plot Generation
The primary objective of this step is to harness the powerful capabilities of GPT-4 to generate diverse and compelling movie plots. Each movie plot comprises basic elements, including overview, characters, and frame descriptions. This aims to construct a high-quality movie plot text for the subsequent style immobilization and generation process. We propose three main strategies for this purpose:
1) We first employ GPT-4 to generate a set of plausible movie scripts. Subsequently, we construct reusable prompts based on these scripts to generate a variety of movie plots in bulk. During this process, phrases describing movie themes, like "a tragic film," serve as variables enabling us to direct GPT-4 in creating movie plots based on specific themes.
2) Within the movie plot text, we incorporate two key elements: character and style. Utilizing the character and style descriptions generated by GPT-4, we aim to ensure great consistency in both the portrayal of characters and the stylistic attributes across key frames in the video.
3) The substantial number of key frames in movie plots poses a challenge, as LLMs are prone to forgetting plot details during continuous generation processes. Therefore, we propose a story expansion strategy to obtain the continuous key frame descriptions and maximally mitigate the forgetting issue inherent in LLMs. We divide the description generation process into three levels.
We now elucidate how we construct the prompt to obtain high-quality movie plot text. Details can be found in appendix B.1. Our prompt is comprised of the following elements:
Movie Theme: As mentioned above, we treat the phrases used to describe a movie theme as variables. Therefore, various movie themes are first generated to ensure the diversity of movie data.
Overview: To provide a referable foundation for the subsequent generation of the movie plot, we initially generate an overview based on the movie theme. This will also facilitate the consistency of the whole movie plot and the generation of summary-type QA.
Style: Another significant element is the movie style which serves as a groundwork for the style inversion process. Appropriate, theme-specific style descriptions can assist in generating scene images related to the movie theme. Through textual inversion, this process facilitates the training of style embeddings. This, in turn, contributes to the consistency of all scenes throughout the movie.
Characters: In movie-level video instruction data generation, ensuring consistency in character portrayal is crucial. Therefore, in the movie plot generation process, we require GPT-4 to play the director role in selecting an appropriate character for movie plots and generating character descriptions. Subsequently, we design to leverage the powerful prior knowledge of celebrity faces from pretrained stable diffusion, aiming to maintain excellent character coherence in the generated key frame portraits.
Story Expanding to Key Frame Description: To guarantee the quantity and quality of key frames, we propose to utilize the aforementioned story expansion strategy, dividing the movie plot into three levels: Epoch Chapters, Narrative Threads, and Frame Description. Each chapter within Epoch Chapters represents a significant period or phase in story development. Narrative Threads further refine each major chapter into more specific narrative strands and sub-stories, showcasing the coherence of character development and main events. Frame Description then elaborates on these narrative strands, depicting the critical visual frames. Then we iteratively expand the story from the preceding level to generate three-level movie plot texts, ultimately yielding key frame descriptions, as illustrated in Fig. 3.
3.2 Style Immobilization Process
The next step is designed to convert the style descriptions from the generated movie plots into information that can guide a T2I model to generate scenes with a consistent style. To achieve this goal, we utilize the previously introduced technique, textual inversion. The specific steps are as follows: 1) Initially, we utilize the stable diffusion to generate arbitrary style scenes based on the specific style descriptions. 2) Subsequently, we employ textual inversion to convert these scene images into embeddings through an optimization step. These embeddings can guide the diffusion model in generating images with specific styles. We refer to this entire process as the style immobilization process, which essentially fixes a certain movie style generated by GPT-4 into the latent space as an embedding, corresponding to a specific token. Thus, in the generation of continuous key frame images, using the specific token solely triggers the diffusion model to generate images in a particular style, ensuring the consistency of scenes across frames.
3.3 Video Instruction Data Generation
After completing the previous two steps, we now have obtained immobilized style embeddings, celebrity-wise characters, and key frame descriptions. Based on these, we utilize the style embeddings to guide stable diffusion in generating key frames according to the key frame descriptions. Furthermore, we progressively generate various instruction QA pairs based on the movie plot. Here are the details.
Consistent Key Frame Generation
It is observed in (Feichtenhofer et al., 2019; Li et al., 2023c; Zhang et al., 2023; Maaz et al., 2023) that key frames at a rather low frame rate have already been able to meet the models’ requirements for understanding video content, thus our method generates key frames based on corresponding descriptions. To generate key frames that are consistent in both characters and scenes, we initially replace character names in the frame description with corresponding celebrities (chosen by GPT-4). Following that, as previously mentioned, a style-immobilized embedding linked to a special token is utilized. This style embedding can serve as a condition to guide the stable diffusion model in generating scenes in a fixed style. This process is triggered by a special token, which, in this paper, denotes the name of a specific style, such as "Gothic" Hence, by incorporating a sentence like "generate an image in Gothic style:" at the beginning of our frame description, combined with celebrity-wise character names, we can generate consistent key frames while maintaining the visual diversity.
QA Pairs Generation
Beyond the textual and visual information inherent to the story itself, enriched QA pairs can well instruct the learning of the model by multimodal information, thereby achieving a superior understanding of movie-level narratives. We utilize GPT-4 to progressively generate rich QA pairs based on all the contents of the movie plot, encompassing aspects of overview, what, where, and why.
@@ 3.3 Video Instruction Data Generation
4 Experiment
In this section, we delineate the experiments conducted to substantiate the effectiveness of our innovative pipeline for data generation tailored to movie-level video instruction tuning. We will first introduce our implementation details and dataset statistics, followed by the evaluation metrics, and finally demonstrate our main results and qualitative results.
4.1 Implementation Details and Dataset Statistics
Implementation Details.
We conduct our video-related experiments on LLaMA-VID. The training procedure of the model is divided into three stages: modality alignment, instruction tuning, and long video tuning. The training conducted during the first two stages endowed the model with the capability to comprehend both images and short videos. Initially, we conducted fine-tuning of LLaMA-VID during the third stage, using their raw long video dataset. We use this trained LLaMA-VID as our base model. Subsequently, we fine-tune the model on our dataset for comparison. Details can be referred to in our appendix.
Dataset Statistics.
We show the statistics of our generated dataset in Table. 1 and Fig. 4. Our instruction tuning data encompasses rich QA pairs that extract in-depth information about video content. We support 15 different genres of movie-level videos, significantly enhancing generalization capability. This is primarily attributed to the robust flexibility of our data construction methodology. Notably, our approach can generate videos of any theme and length, and we plan to continue expanding the richness of our data in the future.
@@ 4.1 Implementation Details and Dataset Statistics
Table 2: Quantitative comparison with other existing methods on quality of generated key frames.
Method MSVD-QA MSRVTT-QA
Acc ↑ Score ↑ Acc ↑ Score ↑
LLaMA-VID 0.493 3.169 0.435 2.865
+Ours 0.567 (+15%) 3.460 (+9%) 0.513 (+18%) 3.141 (+10%)
Table 3: Quantitative comparison with baseline model on two zero-shot video QA datasets.
Metric LLaMA-VID +Ours
Correctness ↑ 1.94 2.154 (+11%)
Detail ↑ 2.431 2.549 (+5%)
Context ↑ 2.701 2.880 (+7%)
Temporal ↑ 1.585 1.832 (+16%)
Consistency ↑ 1.699 1.976 (+16%)
Table 4: Quantitative comparison with baseline model on the video-based generative performance benchmark (Maaz et al., 2023).
Method Overview Plot Temporal
Compare ratio ↑ Score ↑ Compare ratio ↑ Score ↑ Compare ratio ↑ Score ↑
LLaMA-VID 0.39 2.862 0.4 2.657 0.25 1.953
+Ours 0.61 (+56%) 2.881 (+1%) 0.6 (+50%) 2.717 (+2%) 0.75 (+200%) 2.572 (+32%)
Table 5: Quantitative comparison with baseline model on our long video understanding benchmark.

### v2

@@ Abstract
Development of multimodal models has marked a significant step forward in how machines understand videos. These models have shown promise in analyzing short video clips. However, when it comes to longer formats like movies, they often fall short. The main hurdles are the lack of high-quality, diverse video data and the intensive work required to collect or annotate such data. In face of these challenges, we propose MovieLLM, a novel framework designed to synthesize consistent and high-quality video data for instruction tuning. The pipeline is carefully designed to control the style of videos by improving textual inversion technique with powerful text generation capability of GPT-4. As the first framework to do such thing, our approach stands out for its flexibility and scalability, empowering users to create customized movies with only one description. This makes it a superior alternative to traditional data collection methods. Our extensive experiments validate that the data produced by MovieLLM significantly improves the performance of multimodal models in understanding complex video narratives, overcoming the limitations of existing datasets regarding scarcity and bias.
@@
Zhende Song1  Chenchen Wang1  Jiamu Sheng1  Chi Zhang2∗ Gang Yu2  Jiayuan Fan1†  Tao Chen1 1 Fudan University  2 Tencent PCG zdsong23@m.fudan.edu.cn{20307130244, jmsheng18, jyfan, eetchen}@fudan.edu.cn {johnczhang, skicyyu}@tencent.com https://deaddawn.github.io/MovieLLM/
Figure 1: Examples of generated long video instruction data. Our novel approach MovieLLM generate consistent key frames of move-level video with reasonable lines and corresponding question-answer pairs. These data are used to train multimodal large language models on video understanding.
11footnotetext: Project Leader.22footnotetext: Corresponding Author.
@@ 1 Introduction
In recent years, the field of artificial intelligence has witnessed significant progress in the development of large language models (LLMs) such as GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023). These models have shown exceptional ability in generating coherent and contextually relevant text. Further extending the capabilities of LLMs into the multimodal domain, vision language models (VLMs) like LLaVA (Liu et al., 2023a) have demonstrated remarkable proficiency in tasks that require an understanding of both visual and linguistic information, such as image captioning and visual question-answering (VQA). The evolution of VLMs has now begun to include the dynamic realm of video understanding, with models like Video-LLaMA (Zhang et al., 2023a) and VideoChat (Li et al., 2023b) highlighting the potential of VLMs in processing and interpreting video content.
However, a notable gap remains in the ability of these models to understand long-duration videos, such as movies. This limitation is primarily attributed to the lack of extensive long video datasets necessary for tuning these models. LLaMA-VID (Li et al., 2023d) takes a step forward by creating long-video tuning datasets derived from MovieNet (Huang et al., 2020). Nevertheless, constructing such datasets is both cost-intensive and challenging, particularly due to the manual effort required to collect or annotate additional information like movie dialogues. Problems like limited diversity in content can also hinder the generalization capabilities of models.
Motivated by these challenges, we introduce a novel and flexible approach for generating comprehensive datasets for long video instruction tuning. Our methodology leverages the linguistic prowess of GPT-4 and powerful guidance ability of textual inversion technique, offering a unique solution to the problem of dataset diversity and richness. Our pipeline comprises three primary stages:
@@
2) Style Immobilization Process. By adeptly employing textual inversion (Gal et al., 2022), we immobilize the style descriptions generated from the script into the latent space of the diffusion model. This approach guides the model to generate scenes in a consistent style while maintaining diversity under a unified aesthetic.
3) Video instruction data generation. By integrating the powerful generative capabilities of GPT-4 with the developed style-guided diffusion model, we produce style-consistent key frames and corresponding QA pairs, resulting in a comprehensive instruction tuning corpus.
@@
  • We develop a novel pipeline for generating movie-level video instruction tuning datasets by combining GPT-4 and textual inversion.
@@
  • Based on a real movie dataset, we propose a benchmark dataset for evaluating long video comprehension capabilities. Ablation studies on this dataset show the effectiveness of our proposed method significantly boosting model’s understanding on long videos.
@@
2 Related Work
Vision Language Models
With the achievements of large language models (LLMs) such as GPT-4 (OpenAI, 2023) along with their open-source alternatives like LLaMA (Touvron et al., 2023), researchers focus on leveraging the advanced language abilities of LLMs and developing the vision language models (VLMs) that integrate vision models with LLMs for cross-modality understanding. Previous representative VLMs, miniGPT-4 (Zhu et al., 2023a) and LLaVA (Liu et al., 2023a), have shown great capabilities in visual chat by constructing high-quality image-instruction pairs to align the image and text dimensions. Further, VLMs are developed for video understanding (Jin et al., 2024; Li et al., 2023b; Lin et al., 2023; Bai et al., 2024; Ma et al., 2023). Video-LLaMA (Zhang et al., 2023a) utilizes BLIP-2 (Li et al., 2023a) fuse video embedding usig Q-Former. Video-ChatGPT (Maaz et al., 2023) extract video embedding by averaging frame-level features across temporal and spatial dimensions respectively. Besides, some of the recent VLMs focus on 3d information (Li et al., 2023c; Chen et al., 2023a, 2024). However, previous works require generating a large number of tokens for each frame. For hour-long videos, these VLMs are unable to process extensive video sequences. To tackle this problem, LLaMA-VID (Li et al., 2023d) is proposed for long video understanding by encoding each frame with only two tokens.
Video Instruction Tuning Datasets
Preparing video-text pairs to construct video instruction tuning datasets is pivotal in facilitating the training of VLMs for video understanding. Most existing methods (Maaz et al., 2023; Tang et al., 2023; Zhong et al., 2022; Gao et al., 2021; Castro et al., 2022; Yang et al., 2023) construct a video instruction tuning dataset based on benchmark datasets, utilizing rich annotation information. However, most existing methods lack long video data in the constructed instruction tuning datasets, resulting in poor long-video understanding capabilities of the trained models. LLaMA-VID addresses this by extracting data from the MovieNet dataset (Huang et al., 2020) to build an instruction dataset containing movie-level long videos. However, such datasets are limited by the types of annotations, leading to insufficient diversity in the data. Leveraging our novel pipeline, we propose a diverse movie-level video instruction tuning dataset for enhancing long video understanding.
Long Video Generation
Some of the existing works on long video generation utilize either an autoregressive structure or a parallel structure. Those with autoregressive structure (Villegas et al., 2023; Wang et al., 2023; Qiu et al., 2023; Chen et al., 2023b) iteratively apply short video generation models using a sliding window strategy. These methods usually suffer from severe inconsistency. (Yin et al., 2023) use a parallel structure which needs massive well-annotated data for training. Neither of the two considers generating related QA pairs as tuning data. The data they generate is mostly confined to the image styles from training set, allowing for limited customized choices. Moreover, the films they produce predominantly utilize simplistic textual information and lack complex plots and scenes. Other works (Zhu et al., 2023b) use LLM to first generate scripts and then tune a pretrained stable diffusion to generate visuals. However, their data generation domain similarly lies in the training set. Distinct from previous works, our MovieLLM, integrating LLM with textual inversion, can autonomously generate a rich and diverse set of movie data along with corresponding dialogues and QA pairs. It effectively addresses the issue of content inconsistency and presents a superior alternative to traditional data collection methods.
@@ 2 Related Work
3 Method
In this section, we introduce our novel approach for generating movie-level video instruction tuning data. Our method compromises three interrelated steps: movie plot generation, style immobilization process, and video instruction data generation. The complete process is depicted in Fig. 2. The details of these steps are provided in the following subsections.
3.1 Movie Plot Generation
The primary objective of this step is to harness the powerful capabilities of GPT-4 to generate diverse and compelling movie plots. Each movie plot comprises elements including overview, movie style, characters and frame descriptions, which also serves for latter process. For elements like overview, movie style and characters, simple prompt engineering will crack it.
However, our ultimate goal is to generate consistent key frames. To achieve this, a detailed description is needed for each frame. This task is affected by the memory limitations of LLMs. Generating all key frame descriptions simultaneously often results in insufficient context, leading to inconsistent and poor-quality text. To address this issue, we introduce story planner, a strategy using LLM to progressively expand the overview of the story into detailed frame descriptions. This expansion process is divided into three levels. Given the overview $O$ and instruction $I$, LLM $G$ is used to predict level-1 descriptions:
$C_{1},C_{2},...,C_{i},...,C_{l},Q_{L1}=G\left(O,I\right)$ (1)
where $C_{i}$ denotes the current level-1 description, $Q_{L1}$ denotes the corresponding QA pairs for level-1 and $l$ denotes the required length of the descriptions.
Based on $C_{i}$, level-2 descriptions are then further predicted:
$S_{i,1},S_{i,2},...,S_{i,j},...,S_{i,m},Q_{L2}=G\left(C_{i},O,I\right)$
where $S_{i,j}$ denotes the current level-2 description, $Q_{L2}$ denotes the corresponding QA pairs for level-2 and $m$ denotes the required length of the descriptions.
Finally, based on $C_{i}$ and $S_{i,j}$, desired frame descriptions are predicted:
$F_{i,j,1},F_{i,j,2},...,F_{i,j,k},...,F_{i,j,n},Q_{L3},D=G\left(S_{i,j},C_{i},O,I\right)$
where $F_{i,j,k}$ denotes the current frame description, $Q_{L3}$ denotes the corresponding QA pairs for level-3, $D$ denotes the corresponding dialogues and $n$ denotes the required length of the frame descriptions.
We then get totally $l\times m\times n$ frame descriptions and $len(Q_{L1})+len(Q_{L2})+len(Q_{L3})$ QA pairs.
3.2 Style Immobilization Process
Now the issue lays in how to convert text information from the first stage into corresponding visuals. Simply employing a T2I model like stable diffusion for image generation does not guarantee consistency between frames. The key to maintaining consistency lies in ensuring stylistic uniformity throughout the movie. To achieve this, techniques like DreamBooth are generally introduced. However, these methods usually need fine-tuning the entire model, which is clearly impractical when generating movies in bulk. Therefore, in light of these issues, we propose a method based on textual inversion to satisfy the requirements of both consistency and efficiency.
Textual inversion is typically employed in conjunction with Latent Diffusion Models (LDMs). As is known, the goal of LDMs is to remove the noise added to a latent representation of an image. The LDM loss is given by:
$L_{LDM}:=\mathbb{E}_{z\sim\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(z_{t},t,c_{\theta}(y))\|_{2}^{2}\Big{]}\,$
where $t$ is the time step, $z_{t}$ is the latent noised, $\epsilon$ is the unscaled noise sample, $\epsilon_{\theta}$ is the denoising network, $c_{\theta}$ is the encoder for conditioning, $y$ is the conditioning input.
Assume $c_{\theta}$ is a text encoder, text input $y$ is first converted to tokens, which are indexes in some pre-defined dictionary. Each token is then linked to a unique embedding vector that can be obtained through an index-based lookup. These embedding vectors are typically learned as part of the text encoder $c_{\theta}$. The embedding space is the optimized objective of textual inversion. Now, in the first stage of MovieLLM, we generate a word that describes the style of the movie like “Gothic”. We use this word to represent the style concept we wish to learn. The learned embedding vector $V_{*}$ will be used to replace the associated vector of the word like “Gothic”. Then we can include the word that describes the style in sentences like “Generate an image in Gothic style…”. In this way, consistency between frames can be assured.
To obtain the learned embedding as mentioned, a set of images (we use 10) which depicts the target style are needed. We generate those images using the movie style from stage one. The embedding related to the style is then learned through direct optimization, by minimizing the LDM loss of Equation (4) over images from the set. The optimization goal can be defined as:
$v_{*}=\operatorname*{arg\,min}_{v}\mathbb{E}_{z\sim\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(z_{t},t,c_{\theta}(y))\|_{2}^{2}\Big{]}$
while $c_{\theta}$ and $\epsilon_{\theta}$ is fixed. This is a reconstruction task that motivate the learned embedding to capture fine visual details unique to the style concept.
Now, by our method, every movie can have their own style-related embedding through just minutes training.
3.3 Video Instruction Data Generation
Based on previous two stages, we utilize the style embeddings to guide stable diffusion in generating key frames according to the key frame descriptions. More specifically, to generate key frames that are consistent in both characters and scenes, we initially replace character names in the frame description with corresponding celebrities (chosen by GPT-4). Following that, a style-immobilized embedding linked to a special token is utilized. This style embedding can serve as a condition to guide the stable diffusion model in generating scenes in a fixed style. This process is triggered by a special token, which denotes a specific style, such as “Gothic”. Hence, by incorporating a sentence like “generate an image in Gothic style:” at the beginning of our frame description, combined with character names, we can generate consistent key frames while maintaining the visual diversity. We construct around 150 movies with approximately 1.5k frames for each one.
Figure 4: Comparison of the key frames generated by ours and other methods. From left to right is the sequence of events in the story. Our method generates more consistent and text-aligned frames. For example, the fifth event mentions "through the crowd". Both Custom Diffusion and Cones-V2 fail to generate the "crowd".
@@ 3.3 Video Instruction Data Generation
Table 2: Quantitative comparison with other existing methods on quality of generated key frames. Our method attains best performance across all the metrics.
Method MSVD-QA MSRVTT-QA ActivityNet-QA
Acc Score Acc Score Acc Score
VideoLLaMA(Zhang et al., 2023a) 51.6 2.5 29.6 1.8 12.4 1.1
LLaMA-Adapter(Zhang et al., 2023b) 54.9 3.1 43.8 2.7 34.2 2.7
VideoChat(Li et al., 2023b) 56.3 2.8 45.0 2.5 26.5 2.2
Video-ChatGPT(Maaz et al., 2023) 64.9 3.3 49.3 2.8 35.2 2.7
MovieLLM 63.2 3.5 52.1 3.1 43.3 3.3
Table 3: Comparison with leading methods on zero-shot video QA datasets. MovieLLM means training LLaMA-VID with synthetic data from MovieLLM. Utilizing a substantial amount of synthetic tuning data, our method attains superior performance on almost all the metircs and come close to Video-ChatGPT in terms of accuracy on MSVD-QA.
Method Correctness Detail Context Temporal Consistency
VideoLLaMA(Zhang et al., 2023a) 1.96 2.18 2.16 1.82 1.79
LLaMA-Adapter(Zhang et al., 2023b) 2.03 2.32 2.30 1.98 2.15
VideoChat(Li et al., 2023b) 2.23 2.50 2.53 1.94 2.24
Video-ChatGPT(Maaz et al., 2023) 2.40 2.52 2.62 1.98 2.37
MovieLLM 2.64 2.61 2.92 2.03 2.43
Table 4: Comparison with leading methods on video-based generative performance benchmark. MovieLLM means training LLaMA-VID with synthetic data from MovieLLM. Our method attains best performance across all the metircs.
4 Experiment
In this section, we report experiments conducted to validate the effectiveness of dataset generated by MovieLLM tailored to movie-level video instruction tuning.
4.1 Experimental Setup
Implementation Details.
For key frames quality comparison, we compare MovieLLM with Custom diffusion (Kumari et al., 2023) and Cones-V2 (Liu et al., 2023b) by following their original settings. For visual understanding and ablation studies, we train our VLM model based on LLaMA-VID which provides basic abilities crucial for video understanding. Specifically, we conduct tuning after the third stage of LLaMA-VID. For all video inputs, we extract frames at a speed of 1 FPS at 224 resolution. More details can be referred to in our appendix.
Datasets
In this work, we primarily construct the training dataset from our pipeline and MovieNet. Particularly, there are totally 425 movies, with 325 from MovieNet and 100 from ours. We divide the MovieNet dataset into three parts to meet the requirements of instruction tuning and ablation studies. The first part contains 125 movies (part 1, corresponding to approximately 2.6K QA pairs), the second part contains 100 movies (part 2, corresponding to approximately 2.1K QA pairs), and the third part contains 100 movies (part 3). Our 100-movie dataset also corresponds to approximately 2.1K QA pairs (part 4), same with the part 2. We use part 1 as long video pre-training dataset for the third stage of LLaMA-VID, and the model trained in this way serves as our base model. Part 2 and part 4 are used as the tuning datasets for experiments. For part 3, we employ a hybrid approach of human and GPT annotation (80% human, 20% GPT) and utilize it as the evaluation dataset for long videos. As for other evaluation datasets, we use MSVD (Chen and Dolan, 2011), MSRVTT (Xu et al., 2016), ActivityNet (Caba Heilbron et al., 2015), the benchmark from Video-ChatGPT (Maaz et al., 2023), and Ego4D (Grauman et al., 2021; Bärmann and Waibel, 2022).
Dataset Ratio MSVD-QA MSRVTT-QA ActivityNet-QA
Acc Score Acc Score Acc Score
baseline 0 49.3 3.2 43.5 2.9 40.9 3.1
+0.2K 0.1 50.1 3.1 44.2 2.8 40.1 3.0
+1.1K 0.5 52.8 3.3 46.8 2.9 41.5 3.1
+1.7K 0.8 57.6 3.3 49.3 3.1 42.6 3.2
+2.1K 1 63.2 3.5 52.1 3.1 43.3 3.3
Table 5: Ablation study for the impact of our data propotion on zero-shot video QA datasets. Through this ablation study, we can see the impact of varying proportions of our dataset on model performance. With the increase of data proportions, model performance improves accordingly. This indicates the effectiveness of our generated data
Dataset MSVD-QA MSRVTT-QA ActivityNet-QA
MovieNet-1 (2.6K) MovieNet-2 (2.1K) MovieLLM (2.1K) Acc Score Acc Score Acc Score
49.3 3.2 43.5 2.9 40.9 3.1
54.8 3.3 50.1 3.1 42.4 3.1
63.2 3.5 52.1 3.1 43.3 3.3
Table 6: Ablation study for comparison between data from MovieLLM and MovieNet on zero-shot video QA datasets. The results indicate that tuning with our dataset brings an obvious performance upgradation.
Dataset Correctness Detail Context Temporal Consistency
MovieNet-1 (2.6K) MovieNet-2 (2.1K) MovieLLM (2.1K)
1.94 2.43 2.70 1.59 1.70
2.17 2.52 2.56 1.95 1.91
2.64 2.61 2.92 2.03 2.43
Table 7: Ablation study for comparison between data from MovieLLM and MovieNet on video-based generative performance benchmark. The outcomes show that tuning with our dataset leads to improved performance across all the metrics
Dataset BLEU METEOR ROUGE
MovieNet-1 (2.6K) MovieNet-2 (2.1K) MovieLLM (2.1K)
54.85 12.77 13.26
57.66 12.79 13.48
60.24 13.55 13.54
Table 8: Ablation study for comparison between data from MovieLLM and MovieNet on QAEgo4d. The results on standard machine translation metrics show that our dataset can effectively enhance model understanding on long video understanding.
Dataset Overview Plot Temporal
MovieNet-1 (2.6K) MovieNet-2 (2.1K) MovieLLM (2.1K)
2.86 2.66 1.95
2.91 2.87 2.20
3.15 2.93 2.31
Table 9: Ablation study for comparison between data from MovieLLM and MovieNet on Test dataset we construct. The results show that tuning with our dataset can effectively enhance model understanding on several aspects of long videos.